# CleanAgent: Automating Data Standardization with LLM-based Agents

### v3

@@ Abstract.
Data standardization is a crucial part of the data science life cycle. While tools like Pandas offer robust functionalities, their complexity and the manual effort required for customizing code to diverse column types pose significant challenges. Although large language models (LLMs) like ChatGPT have shown promise in automating this process through natural language understanding and code generation, it still demands expert-level programming knowledge and continuous interaction for prompt refinement. To solve these challenges, our key idea is to propose a Python library with declarative, unified APIs for standardizing different column types, simplifying the LLM’s code generation with concise API calls. We first propose Dataprep.Clean, a component of the Dataprep Python Library, significantly reduces the coding complexity by enabling the standardization of specific column types with a single line of code. Then, we introduce the CleanAgent framework integrating Dataprep.Clean and LLM-based agents to automate the data standardization process. With CleanAgent , data scientists only need to provide their requirements once, allowing for a hands-free process. To demonstrate the practical utility of CleanAgent , we developed a user-friendly web application, allowing SIGMOD attendees to interact with it using real-world datasets.
@@ 1. Introduction
We also built a web interface for CleanAgent . It allows the SIGMOD attendees to choose sample data and communicate with CleanAgent for standardization. We provide the demonstration video, which can be found on Youtube.
@@ 2. Type-Specific Standardization API Design
where clean_type is the function name, type represents the type of the current column. The first argument df represents the input DataFrame, the second argument column_name is the column being standardized, and the third argument target_format is the target standardization format users specified. Our API design is flexible and extensible, which makes it convenient for users to add their standardization functions for new data types. Currently, we have 142 standardization functions in Dataprep.Clean, each handles one data type.
@@ 3. CleanAgent Workflow
Basic Structure of LLM-based Agents. According to the previous surveys on LLM-based Agent (Xi et al., 2023), we conclude the basic structure of LLM-based Agents shown in Figure 2. An LLM-based agent includes four main components: (1) a backbone LLM used to generate replies for input prompts, (2) a memory used to store historical conversation messages, (3) a system message defining the role of the agent, and (4) a set of external tools which can be called by the LLM-based agent to complete specific tasks, such as web searching, code execution, etc.
@@
Detailed Workflow. The detailed workflow of CleanAgent is shown in Figure 3. The CleanAgent is composed of four agents, including a Chat Manager, a Column-type Annotator, a Python Programmer, and a Code Executor. They can communicate with each other and automatically complete the data standardization process by cooperation. As mentioned in Figure 2, each agent has its own memory to store the historical conversational messages between it and other agents. Note that the memory of the Chat Manager is uniquely comprehensive, encompassing the entire historical conversational messages from all agents within the CleanAgent system. This extensive memory enables every agent in the CleanAgent to generate responses that are informed by the complete historical messages.
@@
The input of CleanAgent includes a table that needs to be standardized. Data scientists can also input extra requirements such as “the format of the date type column should be MM/DD/YYYY”. By receiving the input table and data scientists’ extra requirements, CleanAgent stores this information into the Chat Manager’s memory and then completes the data standardization process. The Chat Manager delivers messages in its memory to the Column-type Annotator(① in Figure 3). Then, The Column-type Annotator receives the table information and leverages an LLM to annotate the type of each column in the input table. If the The Column-type Annotator cannot figure out the specific type of one column, the Column-type Annotator outputs “I do not know”. The annotation result is returned to the Chat Manager and stored in the Chat Manager’s memory (② in Figure 3). Thirdly, the Python Programmer receives historical messages from the Chat Manager including the column-type annotation results (③ in Figure 3), picks up the corresponding clean functions, and generates Python code for the data standardization process. The generated Python code is also returned to the Chat Manager and stored in the Chat Manager’s memory (④ in Figure 3). Finally, the Code Executor receives historical messages from the Chat Manager including the column-type annotation results and the generated Python code (⑤ in Figure 3), then executes the generated Python code. If the generated code executes without errors, the standardized table is returned; otherwise, the error message is returned to the Chat Manager and stored in its memory (⑥ in Figure 3). Then, CleanAgent will retry the whole workflow until it can complete the data standardization process successfully.
4. Demonstration Scenarios
The demonstration setup requires a laptop that can connect to the Internet so the audience can use CleanAgent smoothly with the LLM API services (we used OpenAI’s GPT). If the conference Internet fails, a mobile hotspot (established via cell phone) can also be used for running CleanAgent .
@@ 5. Conclusion
In this demo paper, we proposed CleanAgent to automate the data standardization process with Dataprep.Clean and LLM-based Agents. We implemented CleanAgent as a web application to visualize the conversations among agents. Other tasks in the data science life cycle, such as data cleaning and data visualization, can also be completed by LLM-based agents (Xue et al., 2023). In the future, it is promising that the data science life cycle can be automatically planned and completed by LLM-based agents’ cooperation.

### v4

@@ Abstract.
Data standardization is a crucial part of the data science life cycle. While tools like Pandas offer robust functionalities, their complexity and the manual effort required for customizing code to diverse column types pose significant challenges. Although large language models (LLMs) like ChatGPT have shown promise in automating this process through natural language understanding and code generation, it still demands expert-level programming knowledge and continuous interaction for prompt refinement. To solve these challenges, our key idea is to propose a Python library with declarative, unified APIs for standardizing different column types, simplifying the LLM’s code generation with concise API calls. We first propose Dataprep.Clean, a component of the Dataprep Python Library, significantly reduces the coding complexity by enabling the standardization of specific column types with a single line of code. Then, we introduce the CleanAgent framework integrating Dataprep.Clean and LLM-based agents to automate the data standardization process. With CleanAgent , data scientists only need to provide their requirements once, allowing for a hands-free process.
@@ 1. Introduction
We also built a web interface for CleanAgent . It allows the users to choose sample data and communicate with CleanAgent for standardization. We provide the demonstration video, which can be found on Youtube.
@@ 2. Type-Specific Standardization API Design
where clean_type is the function name, type represents the type of the current column. The first argument df represents the input DataFrame, the second argument column_name is the column being standardized, and the third argument target_format is the target standardization format users specified. Our API design is flexible and extensible, which makes it convenient for users to add their standardization functions for new data types. Currently, we have 142 standardization functions in Dataprep.Clean, each handles one data type. These functions serve to demonstrate the value of a more declarative approach, illustrating that building declarative data standardization tools for LLMs is not only feasible but essential, motivating the community to develop even more advanced tools.
@@ 3. CleanAgent Workflow
Basic Structure of LLM-based Agents. According to the previous surveys on LLM-based Agent (Xi et al., 2023), an LLM-based agent includes four main components: (1) a backbone LLM used to generate replies for input prompts, (2) a memory used to store historical conversation messages, (3) a system message defining the role of the agent, and (4) a set of external tools which can be called by the LLM-based agent to complete specific tasks, such as web searching, code execution, etc.
@@
Detailed Workflow. The detailed workflow of CleanAgent is shown in Figure 2. The CleanAgent is composed of four agents, including a Chat Manager, a Column-type Annotator, a Python Programmer, and a Code Executor. They can communicate with each other and automatically complete the data standardization process by cooperation. Each agent has its own memory to store the historical conversational messages between it and other agents. Note that the memory of the Chat Manager is uniquely comprehensive, encompassing the entire historical conversational messages from all agents within the CleanAgent system. This extensive memory enables every agent in the CleanAgent to generate responses that are informed by the complete historical messages.
@@
The input of CleanAgent includes a table that needs to be standardized. Data scientists can also input extra requirements such as “the format of the date type column should be MM/DD/YYYY”. By receiving the input table and data scientists’ extra requirements, CleanAgent stores this information into the Chat Manager’s memory and then completes the data standardization process. The Chat Manager delivers messages in its memory to the Column-type Annotator(① in Figure 2). Then, The Column-type Annotator receives the table information and leverages an LLM to annotate the type of each column in the input table. If the The Column-type Annotator cannot figure out the specific type of one column, the Column-type Annotator outputs “I do not know”. The annotation result is returned to the Chat Manager and stored in the Chat Manager’s memory (② in Figure 2).
Thirdly, the Python Programmer receives historical messages from the Chat Manager including the column-type annotation results (③ in Figure 2), picks up the corresponding clean functions, and generates Python code for the data standardization process. The generated Python code is also returned to the Chat Manager and stored in the Chat Manager’s memory (④ in Figure 2). Finally, the Code Executor receives historical messages from the Chat Manager including the column-type annotation results and the generated Python code (⑤ in Figure 2), then executes the generated Python code. If the generated code executes without errors, the standardized table is returned; otherwise, the error message is returned to the Chat Manager and stored in its memory (⑥ in Figure 2). Then, CleanAgent will retry the whole workflow until it can complete the data standardization process successfully.
4. Experiments
Dataset. In our experiment, we employ the Flights dataset from (Rekatsinas et al., 2017), as it contains highly irregular datetime formats across four attributes: scheduled_dept, actual_dept, scheduled_arrival, and actual_arrival. The datetime values in these columns exhibit a wide variety of inconsistent formats, such as "2011-12-08 3:50:00 PM", "2:30pDec 27", "06:45 AM Sun 25-Dec-2011", etc. This makes the dataset particularly suitable for evaluating the standardization capabilities of different systems.
Baselines. We compare CleanAgent with the following two baselines: (1) GPT-4o + Prompting. Data standardization code can be directly generated by prompting powerful chat models such as GPT-4o. (2) Cocoon (Zhang et al., 2024). Cocoon is a one-shot data cleaning system that decomposes complex cleaning tasks into manageable components within a workflow designed to mimic human cleaning processes, leveraging large language models. It supports a variety of data cleaning tasks, including missing value imputation, outlier detection, and functional dependency violation. In this paper, however, we focus on evaluating Cocoon’s ability for data standardization.
Note that there are other LLM-based data cleaning approaches, such as RetClean (Eltabakh et al., 2024). However, RetClean primarily adopts a retrieval-based strategy such as RAG to enhance the ability of LLMs for data cleaning, which supplements the LLM with user-provided data sources. This paradigm is not suitable for our scenario.
Ground Truth Generation. We find that GPT-4o can reliably convert individual datetime strings into a target format (e.g., YYYY-MM-DD HH:MM:SS). Thus, we use it to generate cell-level ground truth values and compile them into a complete table.
Metrics. We use the average cell-level matching rate across all columns as our evaluation metric. For a given table , the cell-level matching rate is computed as:
where is the indicator function, and and denote the standardized and ground truth tables, respectively.
Implementation. Implementation. CleanAgent is implemented in Python 3.10.6. Cocoon is run using its official Colab notebook111https://colab.research.google.com/github/Cocoon-Data-Transformation/cocoon/blob/main/demo/Cocoon_Stage_Demo.ipynb from the GitHub repo222https://cocoon-data-transformation.github.io/page/clean. All methods use the gpt-4o-2024-08-06 model. Experiments are conducted on a MacBook Pro with an M1 chip, 16GB RAM, running macOS Sequoia 15.5.
Results. Table 1 presents the comparison of different systems in terms of cell-level matching rate and latency. CleanAgent achieves a 42.5% cell-level matching rate, approximately 2 higher than that of GPT-4o and Cocoon. These results demonstrate that CleanAgent ’s type-specific standardization API enhances the LLM’s ability to generate more precise and concise standardization code. In addition to higher accuracy, CleanAgent also exhibits lower latency compared to Cocoon. This is because Cocoon generates a one-shot SQL query for all columns without the ability to target specific ones, leading to unnecessary overhead.
System | Cell-Level Matching Rate(%) | Latency (s) |
GPT-4o | 22.0 | 19.76 |
Cocoon | 21.5 | 636.62 |
CleanAgent | 42.5 | 29.57 |
5. User Interface of CleanAgent
We developed a web-based user interface for CleanAgent , allowing users to simply upload their tables without performing any operations. The system then automatically returns the standardized results of their data.
@@ 6. Conclusion
In this paper, we proposed CleanAgent to automate the data standardization process with Dataprep.Clean and LLM-based Agents. We implemented CleanAgent as a web application to visualize the conversations among agents. Other tasks in the data science life cycle, such as data cleaning and data visualization, can also be completed by LLM-based agents (Xue et al., 2023). In the future, it is promising that the data science life cycle can be automatically planned and completed by LLM-based agents’ cooperation.
Appendix A Prompts of Component in CleanAgent
Appendix B Detailed Experiment Settings
B.1. Prompt of GPT-4o Baseline
B.2. Detailed GPT Settings
For CleanAgent , we use GPT-4o with a temperature of 0, a timeout of 60 seconds, and a cache seed of 42. For Cocoon (Zhang et al., 2024), we follow the default setting and set the temperature to 1. For the GPT-4o baseline, the temperature is also set to 0 for consistency with CleanAgent .